OpenStudy (anonymous):

Help needed. A debt of $1300 due in 2 years and $1300 due in 5 years is to be repaid by a single payment now. Find out how much the payment is, if the interest rate is 6% compounded monthly.

OpenStudy (anonymous):

do u have the answer so i need to explain, or do you need the answer?

OpenStudy (anonymous):

I need the answer.

OpenStudy (whpalmer4):

\[FV = PV(1+\frac{r}{n})^{nt}\]\[PV = FV(1+\frac{r}{n})^{-nt}\] PV of $1300 due in 2 years at 6% monthly interest: \[PV = 1300(1+\frac{0.06}{12})^{-12*2}=\] PV of $1300 due in 5 years, same terms: \[PV = 1300(1+\frac{0.06}{12})^{-12*5}=\] Those future payments are worth a bit more than $2000 today.